OpenStudy (anonymous):

I am totally confused: (8,9); x+7y=5

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

I do not understand what you mean

OpenStudy (anonymous):

i am trying to figure out how to work this problem. the instructions says write an equation of the line containging the given point and parrell to the given line.

OpenStudy (anonymous):

Write an equation with the point (8,9). Make the equation the same slope as x+7y=5

OpenStudy (anonymous):

can you show me the steps so I can write down because I am totally confused

OpenStudy (anonymous):

Write an equation using point (8,9) in the formula\[(y -y _{1})=m(x -x _{1})\]